, determine the future value at the end of the final year if deposits are made into an account paying annual interest of 12%,Year Cash flow stream $18,300 $15,250 $12,200 $6,100 $3,050
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

a.A=18300*(1.12)^4+15250*(1.12)^3+12200*(1.12)^2+6100*(1.12)+3050

=$75406(Approx).

b.A=18300*(1.12)^5+15250*(1.12)^4+12200*(1.12)^3+6100*(1.12)^2+3050*(1.12)

=$84455(Approx).